OpenStudy (anonymous):

Find the sum of 7a^3+14a+12 and -6a^3+12a^2-7

OpenStudy (anonymous):

How do I solve this? It's a bit different from my last question...

OpenStudy (madhu.mukherjee.946):

just add the like terms together and then say what do you get?

OpenStudy (anonymous):

I had a^3+26a+10 but I know that's not right.

OpenStudy (madhu.mukherjee.946):

7a^3+14a+12 + -6a^3+12a^2-7 =a^3+12a^2+8a+5

OpenStudy (anonymous):

Can you explain how you did that?

OpenStudy (madhu.mukherjee.946):

yes

OpenStudy (madhu.mukherjee.946):

7a^3-6a^3=a^3 {both are like terms}

OpenStudy (madhu.mukherjee.946):

you have only one term 12a^2 and there is no other term with a^2 so,12a^2+0=12a^2

OpenStudy (anonymous):

oh but the - in front of 6 is not for subtracting. it wants me to add -6 to 7

OpenStudy (madhu.mukherjee.946):

14a-6a=8a{again both like terms}

OpenStudy (madhu.mukherjee.946):

no it wants you to subtract because you have a -ve sign in front of 6

OpenStudy (madhu.mukherjee.946):

you gotta check for like terms and see their symbols carefully and then you gotta add or subtract accordingly

OpenStudy (madhu.mukherjee.946):

again 12-5 =7

OpenStudy (anonymous):

"find the sum" so 6 is negative. but 7+ negativ 6 is the same as 7 minus 6 right?

OpenStudy (madhu.mukherjee.946):

yeah

OpenStudy (madhu.mukherjee.946):

anything more you want to know?

OpenStudy (anonymous):

No I'm pretty sure I understand this tyope of question now. I may have another in a second so i'll check that first.

OpenStudy (madhu.mukherjee.946):

okay

OpenStudy (madhu.mukherjee.946):

may i have my medal please:P

OpenStudy (anonymous):

sure how do i do that

OpenStudy (madhu.mukherjee.946):

just click on the best response option

OpenStudy (anonymous):

just to clarify -4x^2 -7x^2 x^2 ------ -10x^2 assuming i'm adding the polynomials?

OpenStudy (madhu.mukherjee.946):

correct

OpenStudy (anonymous):

ok thanks!!!

OpenStudy (madhu.mukherjee.946):

welcome!:)are you new here

OpenStudy (anonymous):

yes i am lol. i have one (hopefully last for now XD) question for you.

OpenStudy (madhu.mukherjee.946):

welcome to openstudy!!

OpenStudy (madhu.mukherjee.946):

say your question

OpenStudy (anonymous):

-4x^2+3xy- y^2 -7x^2- xy +6y^2 x^2+4xy-2y^2 --------------- my answer -10x^2+6xy-(-3y^2)

OpenStudy (anonymous):

i think something is wrong with it...

OpenStudy (madhu.mukherjee.946):

wait

OpenStudy (madhu.mukherjee.946):

nope its partially correct:)

OpenStudy (madhu.mukherjee.946):

-10x^2+6xy+3y^2

OpenStudy (anonymous):

partially? ok what is the incorrect part?

OpenStudy (madhu.mukherjee.946):

that -(-3y^2)

OpenStudy (anonymous):

ok so it is +3y^2

OpenStudy (madhu.mukherjee.946):

what is 6-1-2

OpenStudy (anonymous):

my school program is telling m there should be one more symbol/digit in there somewhere... 6-1-2 is 3

OpenStudy (madhu.mukherjee.946):

good

OpenStudy (anonymous):

ok well i guess it'll tell me what it's looking for when i send in the answers. but the way you showed me does look right.

OpenStudy (madhu.mukherjee.946):

nope this is correct but i don't know if you have posted correctly

OpenStudy (anonymous):

i think i did. my program likes to mess with me haha

OpenStudy (anonymous):

wait i know hat i did XD. I put 6x instead of 6xy. I got it now. Thank you!!!

OpenStudy (madhu.mukherjee.946):

okay:)

OpenStudy (madhu.mukherjee.946):

don't get panicked just keep yourself calm:)

OpenStudy (madhu.mukherjee.946):

before you do maths just sit and breathe in and out deeply from your heart and try to feel your inner core,this will give you immense help and concentration:)